Circular No. 7936

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ea IN NGC 820 T. Puckett, Mountain Town, GA; and J. Newton, Osoyoos, BC, report

the discovery of an apparent supernova (mag 17.7) on an unfiltered CCD frame (limiting mag 18.8) taken with the Puckett Observatory 0.35-m automated patrol telescope on July 21.35 UT. SN 2002ea is located at α = 2^{h}08^{m}25^{s}.08, δ = +14^{o}20′52′′.8 (equinox 2000.0), which is 1′′.3 east and 6′′.2 south of the center of NGC 820. The new object was confirmed on CCD frames taken on July 22.38 by Puckett with a 0.60-m reflector, but it does not appear on unfiltered images taken by Puckett on 2001 Nov. 11, 22, 2002 Jan. 13, and Feb. 2 (limiting mag ∼ 20.5) or on Palomar Sky Survey images taken on 1991 Oct. 12, 1990 Nov. 18 (limiting mag ∼ 21.0), and 1950 Aug. 14 (limiting mag ∼ 19.7).

COMETS C/2001 C7, 2001 D1, 2002 N1, 2002 O1, 2002 O2, 2002 O3

Further to IAUC 7935, D. Hammer reports measurements for six additional comets found on SOHO website images by S. Hoenig (C/2002 O2; C3 coronagraph), by R. Kracht (C/2001 C7 and C/2001 D1; C2 coronagraph), by X. Leprette (C/2002 N1; C3 coronagraph), and by R. Matson (C/2002 O1 and C/2002 O3; C3 coronagraph). The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below. C/2001 C7 belongs to the Meyer group (cf. IAUC 7832); C/2002 N1, C/2002 O1, C/2002 O2, and C/2002 O3 are Kreutz sungrazers.

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 C7 Feb. 11.579  21^{h}45.^{m}6  −12^{o}18′  2002-O27
C/2001 D1  18.406  21 59.7  −11 31  2002-O27
Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 N1 July 8.431  6^{h}58.^{m}4  +20^{o}40′  2002-O27
C/2002 O1  16.596  7 26.1  +18 17  2002-O33
C/2002 O2  16.596  7 25.5  +18 16  2002-O33
C/2002 O3  16.929  7 31.7  +19 20  2002-O33

COMET 57P/DU TOIT-NEUJMIN-DELPORTE Corrigendum. On IAUC 7935, line 7, for S read T

2002 July 22 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother